OpenStudy (anonymous):

A certain sum is invested for a certain time. It amounts to Rs 400 if the rate is 6% per annum. But when invested at a rateof 2% per annum, it amounts to Rs 250. What is the sum?

OpenStudy (shrutipande9):

see take the principle as x. so in case 1 ur intrest becomes 400-x use the formula I=PNR

OpenStudy (shrutipande9):

show me d equation....

OpenStudy (shrutipande9):

ill be right back...

OpenStudy (anonymous):

A certain sum is invested for a certain time. It amounts to Rs 400 if the rate is 6% per annum. But when invested at a rateof 2% per annum, it amounts to Rs 250. What is the sum?

OpenStudy (shrutipande9):

\[400-x=\frac{ x*N*6 }{ 100 }\]

OpenStudy (shrutipande9):

use this formula and find N first. then instead of 400-x use 250-x and instead of 6 add 2. then find the N. Equate d 2 N and then u will get d ans...

OpenStudy (anonymous):

here N is time ?

OpenStudy (shrutipande9):

yes

OpenStudy (anonymous):

here they mentioned time @shrutipande per annum

OpenStudy (shrutipande9):

see thats for the rate...6% every year...dats not the period for which the principle is kept...

OpenStudy (anonymous):

k i will try

OpenStudy (anonymous):

6Nx=40000-x,25000-x=2Nx

OpenStudy (anonymous):

x=28750

OpenStudy (shrutipande9):

insstead of multipying..keep the equations in this form.\[N=\frac{ 100(400-X) }{ 6X }\] AND \[N=\frac{ 100(250-X) }{ 2X }\]

OpenStudy (shrutipande9):

NOW EQUATE THE N

OpenStudy (anonymous):

its goes lengthier and lengthier

OpenStudy (anonymous):

N=15.33

OpenStudy (shrutipande9):

\[\frac{ 100(400-x) }{ 6x }=\frac{ 100(250-x) }{ 2x }\]

OpenStudy (shrutipande9):

now solve this...

OpenStudy (anonymous):

x=175

OpenStudy (shrutipande9):

cool...u got it:)

OpenStudy (anonymous):

thank u very much

OpenStudy (shrutipande9):

ur welcum:)